OpenStudy (anonymous):

A bomb initially at rest explodes into three equally massive fragments. Right after the explosion, one of the fragments, travels in the positive y-direction with a velocity v1. The two fragments ravel downward, at angles +- 30degrees with respect to the negative y-axis. What is the ratio, R, of the kinetic energy of fragment 1 kinetic energy of fragment 2 (ie. R=k1/k2)

OpenStudy (anonymous):

you might want to use the momentum eq or the mechanical energy conservation.

OpenStudy (anonymous):

Thats what ive been trying. I set the first kinetic energy equal to the sum of the other (.5)m1(v1^2)=(.5)(m2)(v2cos30)^2+(.5)(m3)(v3cos30)^2 All the masses are equal so they can cancell out..but i then i get lost

OpenStudy (anonymous):

Take the velocity of 2nd fragment to be v2 and for 3rd fragment as v3. now apply the conservation of linear momentum in X and Y directions which are as follows - In X direction : \[m v _{2} \sin30 = m v _{3} \sin30 \] which gives V2 = V3 ---------------------------------------------------(1) Now in y direction : \[m v _{1} = m v _{2} \cos30 + m v _{3} \cos 30 \] cancelling out the common terms and putting v2 = v3 we get \[v _{1} = \sqrt{3} \times v _{2} \] Therefore \[K= R _{1} \div R _{2}\] \[= \left( 0.5 \times m v _{1}^{2} \right) \div \left( 0.5 \times m v _{2}^{2} \right)\] and putting \[v _{1} = \sqrt{3} \times v _{2} \] wefinally get K = 3 (Answer)

OpenStudy (anonymous):